# Faster Reinforcement Learning by Freezing Slow States

### v3

@@ Appendix E Proofs for Section 6 and Section 7
Let $\hat{s}=\operatorname*{arg\,max}_{s\in\mathcal{S}}|U_{1}^{*}(s)-U_{2}^{*}(s)|$. We will analyze $\bigl{|}U_{1}^{*}(\hat{s})-U_{2}^{*}(\hat{s})\bigr{|}$.
$\bigl{|}U_{1}^{*}(\hat{s})-U_{2}^{*}(\hat{s})\bigr{|}$ $=\Bigl{|}\max_{a_{1}\in\mathcal{A}}\,\Bigl{\{}r_{1}(\hat{s},a_{1})+\gamma\,\mathbb{E}\bigl{[}U_{1}^{*}\bigl{(}f_{1}(\hat{s},a_{1},w)\bigr{)}\bigr{]}\Bigr{\}}-\max_{a_{2}\in\mathcal{A}}\,\Bigl{\{}r_{2}(\hat{s},a_{2})+\gamma\,\mathbb{E}\bigl{[}U_{2}^{*}(f_{2}(\hat{s},a_{2},w))\bigr{]}\Bigr{\}}\Bigr{|}$
@@
Let state $\hat{s}=\operatorname*{arg\,max}_{s\in\mathcal{S}}\,U_{1}^{*}(\hat{s})-U_{1}^{\tilde{\pi}_{2}}(\hat{s})$ be the state that achieves the largest regret (when using $\tilde{\pi}_{2}$ in $\mathcal{M}_{1}$). Substituting from (29) gives
$U_{1}^{*}(\hat{s})-U_{1}^{\tilde{\pi}_{2}}(\hat{s})$ $=r_{1}(\hat{s},\pi_{1}^{*}(\hat{s}))-r_{1}(\hat{s},\tilde{\pi}_{2}(\hat{s}))+\gamma\,\mathbb{E}\bigl{[}U_{1}^{*}(f_{1}(\hat{s},\pi_{1}^{*}(\hat{s}),w))-U_{1}^{\tilde{\pi}_{2}}(f_{1}(\hat{s},\tilde{\pi}_{2}(\hat{s}),w))\bigr{]}$
@@
$\leq 2\bm{\epsilon}_{r}+2\gamma\bm{\epsilon}_{U}+2\gamma L_{1}(d+\tilde{d})+\gamma\bigl{(}U_{1}^{*}(\hat{s})-U_{1}^{\tilde{\pi}_{2}}(\hat{s})\bigr{)},$
@@
Consider an MDP $\langle\mathcal{S},\mathcal{A},\mathcal{W},f,r,\gamma\rangle$ with reward function $r$ taking values in $[0,r_{\textnormal{max}}]$. Suppose the optimal value function is $U^{*}$ and the associated Bellman operator is $F$. Fix any initial value function such that $0\leq U_{0}(s)\leq r_{\textnormal{max}}/(1-\gamma)$ for all $s$ and let $U^{k}=F^{k}\,U_{0}$ be the result after iteration $k$ of value iteration. Then, it holds that
@@
$&\underbrace{\gamma\,\mathbb{E}\bigl{|}(H^{T-1}U^{*})(x_{1},y_{1})-(\tilde{H}^{T-1}U^{*})(x_{1},y_{1})\bigr{|}}_{\text{TermA}}\\ &+\underbrace{\gamma\,\mathbb{E}\bigl{|}(\tilde{H}^{T-1}U^{*})(x_{1},y_{1})-(\tilde{H}^{T-1}0)(x_{1},y_{1})-\gamma^{T-1}\,U^{*}(x_{T},y_{T})\bigr{|}}_{\text{Term B}},\end{aligned}$
@@
Term $\text{Term A}\;\leq\;\alpha d_{\mathcal{Y}}\biggl{(}L_{r}\sum_{i=1}^{T-2}\gamma^{i}\,\sum_{j=0}^{i-1}L_{f}^{j}+\gamma^{T-1}L_{U}\sum_{j=0}^{T-2}L_{f}^{j}\biggr{)}.$ (30)
@@
where $\mathcal{Y}_{s_{1}}^{T-1}$ is the set of fast states reachable from $s_{1}=(x_{1},y_{1})$ after $T-1$ transitions of $f_{\mathcal{Y}}(x_{1},\cdot,\cdot,\cdot)$. Let $\tilde{y}_{s_{1}}=\operatorname*{arg\,max}_{y\in\mathcal{Y}_{s_{1}}^{T-1}}U^{*}(x_{1},y)$ be the fast state that attains the maximum. Note that $\tilde{y}_{s_{1}}$ depends on $s_{1}$, which is random. Combining with the rest of Term B, we have
@@
where (31) follows by Assumption 2 and (32) comes from Lemma D.1 (we use that $x_{T}(\omega)$ is $T-1$ transitions from $x_{1}(\omega)$ and both $\tilde{y}_{s_{1}}(\omega)$ and $y_{T}(\omega)$ are both $T-1$ transitions from $y_{1}(\omega)$). Finally, we have
[…]
@@
  • For $\bar{U}^{*}$, let $\mathcal{M}_{1}=\langle\mathcal{S},\mathcal{A},\mathcal{W},f_{1},r_{1},\gamma^{T}\rangle$ be the MDP associated with the base model reformulation (8), but with the lower-level policy fixed to be $\bm{\pi}^{*}$. The reward function $r_{1}$ is $r_{1}(s,a)=\mathbb{E}[R(s,a,\bm{\pi}^{*})]$. Given a $T$-period noise sequence $\bm{w}$, an initial state $s$, and action $a$, the “next” state $f_{1}(s,a,\bm{w})=s_{T}(a,\bm{\pi}^{*})$ is the state obtained by first taking action $a$ in state $s$ and then following policy $\bm{\pi}^{*}$ for the next $T-1$ steps.
@@
We also need to account for the fact that $\mu$ is greedy with respect to $V$, an approximation of the optimal value of $\mathcal{M}_{2}$. More precisely, $\mu$ is greedy with respect to $r_{2}(s,a)=\mathbb{E}\bigl{[}\tilde{R}(s_{0},a,J_{1})\bigr{]}$, $f_{2}(s,a,\bm{w})=s_{T}(a,\bm{\pi})$, and value function $V$. We can thus apply Lemma E.2 with $\epsilon_{r}=\epsilon_{r}(\bm{\pi}^{*},J_{1})$, $d=d(\alpha,d_{\mathcal{Y}},T)$, $L_{1}=L_{U}$, $\epsilon_{U}=\bigl{\|}\bar{U}^{*}-V^{*}(J_{1},\bm{\pi})\bigr{\|}_{\infty}$, and $\tilde{\epsilon}_{U}=\bigl{\|}V-V^{*}(J_{1},\tilde{\bm{\pi}})\bigr{\|}_{\infty}$. Collecting terms completes the proof.
@@ Appendix F Bounds on $L_{U}$ in Terms of $L_{r}$ and $L_{f}$
Define the state-action value function $Q(s,a)=r(s,a)+\gamma\,\mathbb{E}\bigl{[}U(f(s,a,w))\bigr{]}$. Then, for any state-action pairs $(s,a)$ and $(\tilde{s},\tilde{a})$, the state-action value function $Q$ satisfies
@@
$\bigl{|}Q(s,a)-Q(\tilde{s},\tilde{a})\bigr{|}$ $\leq|r(s,a)-r(\tilde{s},\tilde{a})|+\gamma\,\bigl{|}\mathbb{E}\bigl{[}U(f(s,a,w))-U(f(\tilde{s},\tilde{a},w))\bigr{]}\bigr{|}$
@@
Consider an $(\alpha,d_{\mathcal{Y}})$-fast-slow MDP $\langle\mathcal{S},\mathcal{A},\mathcal{W},f,r,\gamma\rangle$. Let $Q:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ be a state-action value function and assume there exists $L_{Q}>0$ where for any states $(s,a)$ and $(\tilde{s},\tilde{a})$,
@@
$\bigl{|}U(s)-U(\tilde{s})\bigr{|}$ $=\bigl{|}\max_{a}\,Q(s,a)-\max_{\tilde{a}}\,Q(\tilde{s},\tilde{a})\bigr{|}$
@@ Appendix G Proofs for Appendix C
Applying the above inequality $T-1$ times yields the first result. Next, we see that

### v4

@@ Appendix E Proofs for Section 6 and Section 7
[…]
@@
$\leq\begin{aligned} \epsilon_{r}&+\gamma\max_{a\in\mathcal{A}}\,\bigl|\mathbb{E}\bigl[U_{1}^{*}\bigl(f_{1}(\hat{s},a,w)\bigr)-U_{1}^{*}\bigl(f_{2}(\hat{s},a,w)\bigr)\bigr]\bigr|\\ &+\gamma\max_{a\in\mathcal{A}}\,\bigl|\mathbb{E}\bigl[U_{1}^{*}\bigl(f_{2}(\hat{s},a,w)\bigr)-U_{2}^{*}\bigl(f_{2}(\hat{s},a,w)\bigr)\bigr]\bigr|\end{aligned}$
@@
$\leq\begin{aligned} 2\bm{\epsilon}_{r}&+2\gamma\bm{\epsilon}_{U}+\gamma\,\mathbb{E}\bigl[U_{1}^{*}(\tilde{f}_{2}(\hat{s},\tilde{\pi}_{2}(\hat{s}),w))-U_{1}^{*}(\tilde{f}_{2}(\hat{s},\pi_{1}^{*}(\hat{s}),w))\bigr]\\ &+\gamma\,\mathbb{E}\bigl[U_{1}^{*}(f_{1}(\hat{s},\pi_{1}^{*}(\hat{s}),w))-U_{1}^{\tilde{\pi}_{2}}(f_{1}(\hat{s},\tilde{\pi}_{2}(\hat{s}),w))\bigr]\end{aligned}$
$=\begin{aligned} 2\bm{\epsilon}_{r}&+2\gamma\bm{\epsilon}_{U}+\gamma\,\mathbb{E}\bigl[U_{1}^{*}(\tilde{f}_{2}(\hat{s},\tilde{\pi}_{2}(\hat{s}),w))-U_{1}^{*}(f_{1}(\hat{s},\tilde{\pi}_{2}(\hat{s}),w))\bigr]\\ &+\gamma\,\mathbb{E}\bigl[U_{1}^{*}(f_{1}(\hat{s},\pi_{1}^{*}(\hat{s}),w))-U_{1}^{*}(\tilde{f}_{2}(\hat{s},\pi_{1}^{*}(\hat{s}),w))\bigr]\\ &+\gamma\,\mathbb{E}\bigl[U_{1}^{*}(f_{1}(\hat{s},\tilde{\pi}_{2}(\hat{s}),w))-U_{1}^{\tilde{\pi}_{2}}(f_{1}(\hat{s},\tilde{\pi}_{2}(\hat{s}),w))\bigr]\end{aligned}$
[…]
@@
Consider an MDP $\langle\mathcal{S},\mathcal{A},\mathcal{W},f,r,\gamma\rangle$ with reward function $r$ taking values in $0,r_{\textnormal{max}}]$. Suppose the optimal value function is $U^{*}$ and the associated Bellman operator is $F$. Fix any initial value function such that $0\leq U_{0}(s)\leq r_{\textnormal{max}}/(1-\gamma)$ for all $s$ and let $U^{k}=F^{k}\,U_{0}$ be the result after iteration $k$ of value iteration. Then, it holds that
@@
$\leq\begin{aligned} &\underbrace{\gamma\,\mathbb{E}\bigl|(H^{T-1}U^{*})(x_{1},y_{1})-(\tilde{H}^{T-1}U^{*})(x_{1},y_{1})\bigr|}_{\text{Term A}}\\ &+\underbrace{\gamma\,\mathbb{E}\bigl|(\tilde{H}^{T-1}U^{*})(x_{1},y_{1})-(\tilde{H}^{T-1}0)(x_{1},y_{1})-\gamma^{T-1}\,U^{*}(x_{T},y_{T})\bigr|}_{\text{Term B}},\end{aligned}$
@@
$\text{Term A}\;\leq\;\alpha d_{\mathcal{Y}}\biggl(L_{r}\sum_{i=1}^{T-2}\gamma^{i}\,\sum_{j=0}^{i-1}L_{f}^{j}+\gamma^{T-1}L_{U}\sum_{j=0}^{T-2}L_{f}^{j}\biggr).$ (30)
@@
where $\mathcal{Y}_{s_{1}}^{T-1}$ is the set of fast states reachable from $s_{1}=(x_{1},y_{1})$ after T1T-1 transitions of $f_{\mathcal{Y}}(x_{1},\cdot,\cdot,\cdot)$. Let $\tilde{y}_{s_{1}}=\operatorname*{arg\,max}_{y\in\mathcal{Y}_{s_{1}}^{T-1}}U^{*}(x_{1},y)$ be the fast state that attains the maximum. Note that $\tilde{y}_{s_{1}}$ depends on $s_{1}$, which is random. Combining with the rest of Term B, we have
@@
where (31) follows by Assumption 2 and (32) comes from Lemma D.1 (we use that $x_{T}(\omega)$ is T1T-1 transitions from $x_{1}(\omega)$ and both $\tilde{y}_{s_{1}}(\omega)$ and $y_{T}(\omega)$ are both T1T-1 transitions from $y_{1}(\omega)$). Finally, we have
$\text{Terms A}+\text{B}\;$ $\leq\;\alpha d_{\mathcal{Y}}\biggl(L_{r}\sum_{i=1}^{T-2}\gamma^{i}\,\sum_{j=0}^{i-1}L_{f}^{j}+\gamma^{T-1}L_{U}\sum_{j=0}^{T-2}L_{f}^{j}\biggr)+\gamma^{T}L_{U}d_{\mathcal{Y}}(\alpha+2)(T-1)$
@@
  • For $\bar{U}^{*}$, let $\mathcal{M}_{1}=\langle\mathcal{S},\mathcal{A},\mathcal{W},f_{1},r_{1},\gamma^{T}\rangle$ be the MDP associated with the base model reformulation (8), but with the lower-level policy fixed to be $\bm{\pi}^{*}$. The reward function $r_{1}$ is $r_{1}(s,a)=\mathbb{E}[R(s,a,\bm{\pi}^{*})]$. Given a $T$-period noise sequence $\bm{w}$, an initial state $s$, and action $a$, the “next” state $f_{1}(s,a,\bm{w})=s_{T}(a,\bm{\pi}^{*})$ is the state obtained by first taking action $a$ in state $s$ and then following policy $\bm{\pi}^{*}$ for the next T1T-1 steps.
@@
[…]
@@ Appendix F Bounds on $L_{U}$ in Terms of $L_{r}$ and $L_{f}$
[…]
@@
[…]
@@
[…]
@@
[…]
@@ Appendix G Proofs for Appendix C
Applying the above inequality T1T-1 times yields the first result. Next, we see that